OpenStudy (jadenwalker76):

Simplify square root of 5 open parentheses 10 minus 4 square root of 2 close parentheses.

OpenStudy (asylum15):

Jaden, are you here? I can do this for you.

OpenStudy (asylum15):

\[\sqrt{5} (10-4\sqrt{2})\] \[10\sqrt{5} - 4\sqrt{5.2}\] \[10\sqrt{5} - 4\sqrt{10}\] :)